# The Frequency Response of Networks as Open Systems

### v1

@@ II.2 $\mathcal{H}_{2}$-Norm
where $W_{c}$ is the continuous-time controllability Gramian and is defined as
@@
The matrix $W_{c}$ is a symmetric and positive semi-definite matrix and is finite if and only if the matrix $A$ has all eigenvalues with negative real parts. As can be seen from Eq. (4), the $\mathcal{H}_{2}$-norm squared arises when looking both at the response of a system to a given frequency $\omega$ and to a time $t$-signal (note the two integrations over the frequencies $\omega$ and over time $t$); therefore it provides a measure of the amplification when considering arbitrary frequencies between 00 and $\infty$ and arbitrary time signal. The matrix $W_{c}$ can be easily calculated by solving the following continuous-time Lyapunov equation for $W_{c}$:
$AW_{c}+W_{c}A^{\top}+BB^{\top}=0.$ (6)
For more information on the properties of the Controllability Gramian, see Methods IV.2, which in particular discusses the important modularity property of the trace of the output controllability Gramian $\text{Tr}(W_{c}^{out})$. Namely, for any choice of the sets of input nodes and output nodes, the trace $\text{Tr}(W_{c}^{out})$ can be written as a sum of individual contributions, each one of which corresponding to a pair of input and output nodes (see Methods IV.2.) This has direct and important consequences, as it implies that if we assign the set of output nodes, a set of $k$ input nodes that minimizes (maximizes) $\text{Tr}(W_{c}^{out})$ is simply composed of the $k$ nodes that provide the largest (smallest) individual contributions. An analogous conclusion is obtained in the case that a given set of input nodes is assigned and one wants to select an optimal set of $k$ output nodes.
Supplementary Note 1 relates the output standard deviation to the $\mathcal{H}_{2}$-norm via the Gramian spectrum, and Supplementary Note 2 provides a closed-form and bounds for the Bode integral in terms of the eigenvalues and eigenvectors of the matrix $A$.
Up to this point, we have considered the overall network amplification of input signals across different frequencies, as provided by the $\mathcal{H}_{2}$-norm. However, it is possible that a particular network will pass certain frequencies and block others, which can be studied by using Bode plots [42]. To illustrate the response of networks as open systems to individual frequencies, we consider the following example:
Example 1.
Any periodic input signal can be decomposed into different frequencies through its Fourier series decomposition, which allows the application of frequency response analysis to generic periodic input signals. We focus on the unidirectional chain network shown in Fig. 2, where we assume node $1$ is the input node and node $N$ is the output node. For this network, the adjacency matrix $A$ is
$\begin{split}A&=\begin{bmatrix}-c&&&&\\ w_{1}&-c&&&\\ &w_{2}&-c&&\\ &&\ddots&\ddots&\\ &&&w_{N-1}&-c\end{bmatrix}.\end{split}$ (7)
The transfer function for this network is
$G(s)=\dfrac{w_{1}w_{2}\ldots w_{N-1}}{(s+c)^{N}}.$ (8)
The schematics for the Bode plot of this system is shown in Fig. 2 that summarizes our results, where we sketch the Bode plot based on parameters $c,N,w_{1},\ldots,w_{N}$. We see that for frequencies $\omega$ below $c$, the rate of amplification is given by the DC gain, but as $\omega$ grows larger than $c$, the amplitude of the output signal decreases with a rate $\propto 1/\omega^{N}$. The phase plot has a peak and a low plateau corresponding to low and high frequencies, respectively. A negative phase shift arises in the output signal compared to the input signal, by as much as $-90N$ degrees in large frequencies. The cornering frequency is the frequency at which the power output of a system is reduced to half of its passband value. This corresponds to a 3 dB drop in the magnitude of the signal. Fig. 2 shows that the cornering frequency is equal to $c$.
FIG. 2: Bode diagram for chain network. Panel a shows the unidirectional chain network with $N$ nodes and positive link weights $w_{1},w_{2},\ldots,w_{N-1}$ and the self-loops with the weights $-c\leq 0$. Panel b shows the Bode plot corresponding to this system when node $1$ is the input node and node $N$ is the output node. In both plots, the $\omega$-axis is on a log scale.
Equation (2) can be rewritten as follows,
$G(s)=\dfrac{d}{(s+c)^{N}},$ (9)
where $d\in\mathbb{R}$, $c>0$, and $N\in\mathbb{N}$. The $\mathcal{H}_{2}$-norm squared for $G$ is
$\|G\|_{2}^{2}=\dfrac{d^{2}}{2c^{2N-1}}\dfrac{\binom{2N-2}{N-1}}{4^{N-1}}.$ (10)
Table 1 summarizes $\|G\|_{2}^{2}$ for $N=1,2,\ldots,6$. In conclusion, we see that as $N$ grows, the value of $\mathcal{H}_{2}$-norm decreases, which suggests that for larger chains, the dissipation rate $-c$ dominates the amplification rate $d$.
Table 1: The $\mathcal{H}_{2}$-norm squared $\|G\|_{2}^{2}$ in Eq. (10) for different values of the system order $N$.
$N$ 1 2 3 4 5 6
     $\|G\|_{2}^{2}$ $\dfrac{d^{2}}{2c}$ $\dfrac{d^{2}}{4c^{3}}$ $\dfrac{3d^{2}}{16c^{5}}$ $\dfrac{5d^{2}}{32c^{7}}$ $\dfrac{35d^{2}}{265c^{9}}$ $\dfrac{63d^{2}}{512c^{11}}$
II.3 Real Network Analysis
It is natural to expect that many biological networks may have formed throughout evolution in order to achieve a desired balance between passing and blocking behavior. For example, certain networks may require amplifying given input signals, while others may be used to suppress disturbances and noise. At the same time, a crucial question in technological applications is how to modify an existing network to enhance its passing versus blocking characteristics. In this context, both the directedness of the network—reflecting its asymmetry and hierarchical organization—and the functional distinction between input nodes (sources) and output nodes (sinks) are critical in determining signal propagation and amplification. Hereafter, by input or sources we refer to the nodes where energy, matter, or information enter the network from the environment, initiating dynamical activity that propagates through the network. In food webs, for instance, primary producers such as plants or phytoplankton act as inputs by converting external energy (sunlight) into biomass [43, 44, 17]; in power grids, generators inject electrical energy into the network [45, 46, 47, 48, 49, 50, 51]; in connectomes, sensory or afferent regions receive external stimuli and relay signals [52, 53, 54, 55, 56, 57]; in gene regulatory networks, upstream transcription factors respond to environmental cues to regulate downstream targets [58, 59, 17, 60, 61]; and in signaling pathways, receptors at the cell membrane or intracellular sensors detect external signals and activate downstream cascades [62]. These examples underscore that identifying appropriate sources and sinks based on their biological or physical roles is essential for understanding and optimizing the passing or blocking characteristics of real-world networks.
FIG. 3: Real data analysis. Panel a shows the $z$-score vs $p$-value of the real data choice of the input nodes based on evaluation of the trace $\text{Tr}(W_{c})$ of real network data. Panels b-f show the distributions of $\text{Tr}(W_{c})$ for selected real networks over 10,000 sets of randomly chosen input nodes. The number of randomly input nodes is the same as the number of input nodes from the real data. For each network, the choice of real input nodes is plotted as a red star. The violin plots are shifted for better visualization, such that the mean of each distribution lies on the dashed black line. Networks within the same family are plotted using the same color in different panels: b: Pathway networks (blue), c: Food Webs (orange), d: Genetic network (yellow), e: Connectomes (purple), and f: Power Grids (green).
By using the analytical tools introduced earlier, particularly the $\mathcal{H}_{2}$-norm, we now analyze the blocking versus passing behavior of several empirical networks. All networks that we analyze are either unweighted or weighted with non-negative weights (see Supplementary Note 3). For each dataset, we fix the number of input nodes $m$ to its empirical (measured) value. We first choose uniformly at random 10,000 sets of $m$ nodes without repetitions, set them as input nodes and evaluate $\text{Tr}(W_{c})$, the trace of the infinite-horizon continuous-time controllability Gramian. We compare the resulting $\text{Tr}(W_{c})$ from randomly chosen input nodes with the $\text{Tr}(W_{c})$ resulting from input nodes provided by the empirical data. The comparison is performed by evaluating two measures:
[…]
@@ II.3 Real Network Analysis
and we recall that extreme (large magnitude) $z$-scores correspond to low $p$-values, indicating statistically atypical configurations, while small $z$-scores and high $p$-values reflect typical behavior. The results of $z$-score vs $p$-value are shown in Figure 3 a. We consider different categories of networks: power grids for which input nodes are generators, connectomes for which input nodes are sensory neurons, molecular signaling networks for which input nodes are receptors, gene regulatory networks for which input nodes are transcription factors, pathway networks for which input nodes are extracellular ligands, and food webs for which input nodes are autotrophs (plants and algae) that produce their food via photosynthesis. For a detailed explanation of the real network dataset, see Supplementary Note 3. For all networks, in the absence of more detailed information, we set $C=I$. The violin plots in Fig. 3 b-c show the distribution of $\text{Tr}(W_{c})$ along with the real network choice. These distributions and the real network choice are used to evaluate the $z$-score and $p$-value shown in panel a.
@@
Figure 3 shows that pathway networks and food webs, the real choice of input nodes results in passing behavior, while power grids are either blocking or typical. Genetic networks are either passing or typical, and connectomes do not show a discernible pattern since some networks are passing, some are blocking, and the rest are typical. The case of power grids is interesting as these networks are the most blocking. There two reasons for this, the first one being that in these networks, input nodes are the generator buses where the power is injected into the grid, and that the power flows on the transmission lines are given by the voltage angle differences. Typically, the generator buses either have many transmission lines to their neighbors, or a few lines with a high capacity, allowing the power flow to circulate with small angle differences and therefore ensuring the stability of the grid. One thus expect a power grid with well connected generator to be blocking in general. The second reason is that, at odds with the other types of networks considered here, power grids have a matrix $A$ that is almost symmetric. The latter corresponds to the Jacobian of the voltage angle dynamics close to the operational state. In grid where the resistance is negligible compared to the reactance, $A$ is symmetric. When including the transmission line losses, one add a slight asymmetry to $A$ which remains close to a symmetric matrix. The violin plots confirm the same result while providing a more detailed view, as they display the entire distribution: atypical values appear in the tails, whereas typical values cluster near the peak, usually centered around the mean. An explanation for the power grids
@@
Complementary results are reported in Supplementary Notes 4 and 5, where alternative controllability metrics are analyzed. In Supplementary Note 4, we consider the largest eigenvalue of the controllability Gramian, $\lambda_{\max}(W_{c})$, as a measure of the maximum amplification rate achievable by input signals. The findings are analogous to those reported in the main text, confirming the same qualitative distinctions between passing and blocking regimes across network types. Supplementary Note 5 further presents histograms of the $\mathcal{H}_{2}$-norm for empirical networks, comparing two scenarios: (i) all nodes as input nodes and (ii) $m$ randomly selected input nodes, where $m$ equals the number of inputs in the empirical data.
@@ II.4 Directedness as a Passing Mechanism
Systems such as food webs, signaling pathways, and gene regulatory networks tend to exhibit larger $z$-scores and stronger directed hierarchical structure, being almost or strictly DAGs, consistent with their high non-normality and directional flow [36]. Non-normality, which strongly impacts the behavior of dynamical systems [19, 20, 23], is a prominent feature in many naturally occurring networks, as shown in Refs. [18, 36]. In contrast, power grids are either blocking or typical, and neuronal networks do not show a discernible pattern since some networks are passing, some are blocking, and the rest are typical. This is consistent with the fact that power grids tend to be symmetric, in general, and connectomes exhibit low $z$-scores and weak directedness or non-normality (see also Ref. [36]). Based on these structural properties, one can conjecture that most real-world networks, due to their inherent directedness, are expected to exhibit passing behavior under optimal input selection. To shed light on this conjecture, we next examine why the trace of the Gramian (or equivalently the squared $\mathcal{H}_{2}$-norm) varies across network classes, focusing on their directedness as captured by the Henrici index [35].
@@
Figure 4 shows the normalized trace of the controllability Gramian, $\text{Tr}(W_{c})/N^{2}$, versus the Henrici index $\hat{d}_{F}(A)$. We observe a clear trend: higher Henrici index correlates with higher normalized trace. Moreover, networks of the same category cluster together in this plot, revealing distinct levels of non-normality and signal amplification across categories. Notably, food webs, signaling pathways, and gene regulatory networks exhibit a pronounced scale separation—one to two orders of magnitude higher in normalized trace compared to other networks. This is consistent with their (almost) perfect directed acyclic graph (DAG) structure, which maximizes passing behavior along input–output paths. Interestingly, while food webs and pathways achieve this via fewer but longer paths, gene regulatory networks exhibit many very short paths, resembling a collection of directed star graphs [36]. As further detailed in the Supplementary Note 6, the unnormalized $\text{Tr}(W_{c})$ reinforces this observation by clustering gene regulatory networks closer to food webs and pathways, underscoring the role of perfect directedness in enhancing the passing property. We will explore these structural mechanisms more deeply in the next section.
FIG. 4: Normalized trace of Gramian vs. the non-normality index. The normalized Henrici index $\hat{d_{F}}(A)$ is defined in Eq. (12). The dataset is plotted with the same color and marker as in Fig. 3. Pathway, Food web, Genetic, Connectome, and Power Grid networks are in blue, orange, yellow, purple, and green, respectively.

### v2

@@ II.2 $\mathcal{H}_{2}$-Norm
where $w_{i}$ represents the coupling from node $i$ to node $i+1$, i=1,,N1i=1,...,N-1 and $c\leq 0$ is the stabilizing self feedback coefficient at each individual node. The transfer function for this network is [41]
$G(s)=\dfrac{w_{1}w_{2}\ldots w_{N-1}}{(s+c)^{N}}=\dfrac{\xi_{N}}{(s+c)^{N}},$ (8)
where $\xi_{N}=w_{1}w_{2}\dots w_{N-1}$ is the product of all the node-to-node couplings along the chain with $N$ nodes. The schematics for the Bode plot of this system is shown in Fig. 2 that summarizes our results, where we sketch the Bode plot based on the parameters $c,N,w_{1},\ldots,w_{N-1}$. We see that for frequencies $\omega$ below $c$, the rate of amplification is given by the DC gain, but as $\omega$ grows larger than $c$, the amplitude of the output signal decreases with a rate $\propto 1/\omega^{N}$. The phase plot has a peak and a low plateau corresponding to low and high frequencies, respectively. A negative phase shift arises in the output signal compared to the input signal, by as much as 90N-90N degrees in large frequencies. The cornering frequency is the frequency at which the power output of a system is reduced to half of its passband value. This corresponds to a 3 dB drop in the magnitude of the signal. Fig. 2 shows that the cornering frequency is equal to $c$.
FIG. 2: Bode diagram for chain network. Panel a shows the unidirectional chain network with $N$ nodes and positive link weights $w_{1},w_{2},\ldots,w_{N-1}$ and the self-loops with the weights $c\leq 0$. Panel b shows the Bode plot corresponding to this system when node $1$ is the input node and node $N$ is the output node. In both plots, the $\omega$-axis is on a log scale.
The $\mathcal{H}_{2}$-norm squared for $G$ is
$\|G\|_{2}^{2}=\dfrac{\binom{2N-2}{N-1}}{2^{2N-1}}\dfrac{\xi_{N}^{2}}{c^{2N-1}}=\phi_{N}\dfrac{\xi_{N}^{2}}{c^{2N-1}}.$ (9)
The details of the calculation can be found in the Supplementary Material.
Table 1 displays $\phi_{N}$ for $N=1,2,\ldots,6$, from which we see that the coefficients $\phi_{N}$ decay sub-linearly with $N$.
It is important to study the effects of varying the length of the chain $N$ on the $\mathcal{H}_{2}$ norm squared in Eq. (9). We first consider the case that $w_{1}=w_{2}=...=w_{N}=\bar{w}$, then $\xi_{N}=\bar{w}^{N-1}$. For this case, we see that as the length of the chain $N$ grows, $\|G\|_{2}^{2}$ in Eq. (9) will either increase or decrease, depending on whether $|\bar{w}|>c$ or $|\bar{w}|<c$, i.e., whether the magnitude of the node-to-node coupling stimulation exceeds the local suppression at each node. Intuitively, something similar happens in the case of heterogeneous $w_{i}$ couplings, i.e., $\|G\|_{2}^{2}$ will be large (small) when $\tilde{\xi}>c$ ($\tilde{\xi}<c$), where $\tilde{\xi}=|\xi_{N}|^{\frac{1}{N-1}}$ is the geometric mean of the absolute values of the node-to-node couplings along the chain with $N$ nodes.
We also derived a closed form expression for the $\mathcal{H}_{2}$-norm in the case in which the self-feedback coefficients at different nodes along the chain are different $c_{1}\neq c_{2}\neq...\neq c_{N}$,
$\|G\|_{2}^{2}=(-1)^{N}\xi_{N}^{2}\sum_{i=1}^{N}\dfrac{1}{2c_{i}}\prod_{j=1;j\neq i}^{N}\dfrac{1}{(c_{i}^{2}-c_{j}^{2})}\,.$ (10)
See the Supplementary Material Note 1.B for our derivations. Overall, this solvable directed-chain model shows that the chain behaves as a feed-forward filter: when node-to-node stimulation exceeds local damping, signals propagate through the chain, whereas when damping dominates, the chain effectively blocks them. This pass–block behavior mirrors what occurs in more general feed-forward structures in DAG-like networks [18] and, as we demonstrate in the next section, also underpins signal transmission patterns in a large dataset of real-world networks.
Table 1: The coefficient $\phi_{N}$ in Eq. (9) for different values of the system order NN.
NN 1 2 3 4 5 6
$\phi_{N}$ $\dfrac{1}{2}$ $\dfrac{1}{4}$ $\dfrac{3}{16}$ $\dfrac{5}{32}$ $\dfrac{35}{256}$ $\dfrac{63}{512}$
II.3 Analysis of Real Networks
It is natural to expect that many biological networks may have formed throughout evolution in order to achieve a desired balance between passing and blocking behavior. For example, certain networks may require amplifying given input signals, while others may be used to suppress disturbances and noise. At the same time, a crucial question in technological applications is how to modify an existing network to enhance its passing versus blocking characteristics. In this context, both the directedness of the network—reflecting its asymmetry and hierarchical organization—and the functional distinction between input nodes (sources) and output nodes (sinks) are critical in determining signal propagation and amplification. Hereafter, by sources we refer to the nodes where energy, matter, or information enter the network from the environment, initiating dynamical activity that propagates through the network. In food webs, for instance, primary producers such as plants or phytoplankton act as sources by converting external energy (sunlight) into biomass [49, 50, 17]; in power grids, generators inject electrical energy into the network [51, 52, 53, 54, 55, 56, 57]; in connectomes, sensory or afferent regions receive external stimuli and relay signals [58, 59, 60, 61, 62, 63]; in gene regulatory networks, upstream transcription factors respond to environmental cues to regulate downstream targets [64, 65, 17, 66, 67]; and in signaling pathways, receptors at the cell membrane or intracellular sensors detect external signals and activate downstream cascades [68]. These examples underscore that identifying appropriate sources and sinks based on their biological or physical roles is essential for understanding and optimizing the passing or blocking characteristics of real-world networks.
FIG. 3: Real data analysis. Panel a shows the $z$-score vs $p$-value of the real data choice of the input nodes based on evaluation of the trace $\text{Tr}(W_{c})$ of real network data (based on their Jacobian matrix AA and the input matrix $B$). Panels b-f show the distributions of $\text{Tr}(W_{c})$ for selected real networks over 10,000 sets of randomly chosen input nodes. The number of randomly input nodes is the same as the number of input nodes from the real data. For each network, the choice of real input nodes is plotted as a red star. The violin plots are shifted for better visualization, such that the mean of each distribution lies on the dashed black line. Networks within the same family are plotted using the same color in different panels: b: Pathway networks (black), c: Food Webs (red), d: Genetic network (yellow), e: Connectomes (purple), and f: Power Grids (green).
By using the analytical tools introduced earlier, particularly the $\mathcal{H}_{2}$-norm, we now analyze the blocking versus passing behavior of several empirical networks. In this section, to accommodate different types of network dynamics, we consider the general nonlinear model introduced in Sec. A, $\dot{z}_{i}(t)=f(z_{i}(t))+\sigma\sum_{j}\tilde{A}_{ij}\,g(z_{i}(t),z_{j}(t))$, $i=1,..,N$, where the specific forms of the functions $f$ and $g$ depend on the underlying system. These functions vary across applications—for example, they take different forms in food webs, power grids, and connectomes—as detailed in Methods, IV.5.
In the analysis that follows, for each network model, we evaluate the Jacobian $A$ of the nonlinear system about a stable equilibrium, and use the pair $(A,B)$ to calculate the trace of the Gramian $\text{Tr}(W_{c})$ and the $\mathcal{H}_{2}$-norm. These Jacobian matrices may have entries with negative sign. For detailed information on the nonlinear models and the evaluation of the Jacobians, see the Methods IV.5. For each dataset, we fix the number of input nodes $m$ to its empirical (measured) value. We first choose uniformly at random 10,000 sets of $m$ nodes without repetitions, set them as input nodes, and evaluate $\text{Tr}(W_{c})$, the trace of the infinite-horizon continuous-time Gramian. We compare the resulting $\text{Tr}(W_{c})$ from randomly chosen input nodes with the $\text{Tr}(W_{c})$ resulting from input nodes provided by the empirical data. The comparison is performed by evaluating two measures:
$z\text{-score}=\dfrac{x-\bar{x}}{s},\quad p\text{-value}=\int_{x_{\min}}^{x_{real}}p(x)dx\mod 0.5,$ (11)
where $\bar{x}$ is the mean of the sample, $s$ is the standard deviation of the sample, $x_{\min}$ is the minimum of the sample, $x_{real}$ is the real data choice, and $p(x)$ is the probability density function such that $\int_{x_{\min}}^{x_{\max}}p(x)dx=1$. We define
$\text{Passing network}:$ $\begin{cases}\text{Low $p$-value},\\ \\ \text{large positive $z$-score},\end{cases}$
@@ II.3 Analysis of Real Networks
and we recall that extreme (large magnitude) $z$-scores correspond to low $p$-values, indicating statistically atypical configurations, while small $z$-scores and high $p$-values reflect typical behavior. In simple terms, large positive $z$-scores correspond to networks that predominantly pass (amplify) inputs, large negative $z$-scores correspond to networks that predominantly block (attenuate) them, and values near zero indicate networks that are neither particularly passing nor particularly blocking. The results of $z$-score vs $p$-value are shown in Figure 3 a. We consider different categories of networks: power grids for which input nodes are generators, connectomes for which input nodes are sensory neurons, molecular signaling networks for which input nodes are receptors, gene regulatory networks for which input nodes are transcription factors, pathway networks for which input nodes are extracellular ligands, and food webs for which input nodes are autotrophs (plants and algae) that produce their food via photosynthesis. For a detailed explanation of the real network dataset, see Supplementary Note 4. For all networks, in the absence of more detailed information, we set $C=I$. The violin plots in Fig. 3 b-c show the distribution of $\text{Tr}(W_{c})$ along with the real network choice. These distributions and the real network choice are used to evaluate the $z$-score and $p$-value shown in panel a.
@@
Figure 3 shows that in pathway networks, food webs, and genetic networks, the real choice of input nodes results in passing behavior, while power grids and connectomes do not show a discernible pattern, since some networks are passing, some are blocking, and the rest are typical. The case of power grids is interesting as some of these networks are the most blocking. There are two reasons for this, the first one being that in these networks, input nodes are the generator buses where the power is injected into the grid, and the power flows on the transmission lines are given by the voltage angle differences. Typically, the generator buses either have many transmission lines to their neighbors, or a few lines with a high capacity, allowing the power flow to circulate with small angle differences and therefore ensuring the stability of the grid. One thus expects a power grid with well connected generator to be blocking in general. The second reason is that, at odds with the other types of networks considered here, power grids have a Jacobian matrix $A$ that is close to symmetric. The latter corresponds to the Jacobian of the voltage angle dynamics close to the operational state. In a grid where the resistance is negligible compared to the reactance, $A$ is symmetric. Inclusion of the transmission line losses adds a slight asymmetry to $A$, which remains close to a symmetric matrix. The violin plots confirm our conclusions while providing a more detailed view, as they display the entire distribution: atypical values appear in the tails, whereas typical values cluster near the peak, usually centered around the mean. Connectomes represent another class of networks characterized by notably stronger blocking behavior. In Subsection D we explain this in terms of their weak directionality and non-normal properties.
@@
Complementary results are reported in Supplementary Notes 5, where adjacency matrices of the real data is used instead of the Jacobian matrix in the evaluation of the $\mathcal{H}_{2}$-norm, yielding comparable results. In Supplementary Note 6, we consider the largest eigenvalue of the Gramian, $\lambda_{\max}(W_{c})$, as a measure of the maximum amplification rate achievable by input signals. The findings are analogous to those reported in the main text, confirming the same qualitative distinctions between passing and blocking regimes across network types. Supplementary Note 7 further presents histograms of the $\mathcal{H}_{2}$-norm for empirical networks, comparing two scenarios: (i) the case in which all nodes serve as input nodes, and (ii) the case in which $m$ input nodes are selected at random, with $m$ equal to the number of inputs in the empirical dataset.
@@ II.4 Directedness as a Passing Mechanism
Systems such as food webs, signaling pathways, and gene regulatory networks tend to exhibit larger $z$-scores and stronger directed hierarchical structure, being almost or strictly DAGs, consistent with their high non-normality and directional flow [22]. Non-normality, which strongly impacts the behavior of dynamical systems [19, 20, 24], is a prominent feature in many naturally occurring networks, as shown in Refs. [18, 22]. In contrast, power grids and neural networks are either blocking or typical. This is consistent with the fact that power grids tend to be symmetric, in general, and connectomes exhibit weak directedness or non-normality (see also Ref. [22]). Based on these structural properties, one can conjecture that most real-world networks, due to their inherent directedness, are expected to exhibit passing behavior under optimal input selection. To shed light on this conjecture, we next examine why the trace of the Gramian (or equivalently the squared $\mathcal{H}_{2}$-norm) varies across network classes, focusing on their directedness as captured by the Henrici index [39].
@@
Figure 4 shows the normalized trace of the Gramian, $\text{Tr}(W_{c})/N^{2}$, versus
the Henrici index $\hat{d}_{F}(A)$ of the Jacobian matrix $A$. The data is generated in the same way as in Fig. 3. We observe a clear trend: higher Henrici index correlates with higher normalized trace. Moreover, networks in the same category cluster together in this plot, revealing distinct levels of non-normality and signal amplification across categories. Notably, food webs and signaling pathways exhibit a pronounced scale separation—one to two orders of magnitude higher in normalized trace compared to other networks. This is consistent with their (almost) perfect directed acyclic graph (DAG) structure, which maximizes passing behavior along input–output paths. Interestingly, while food webs and pathways achieve this via fewer but longer paths, gene regulatory networks exhibit many very short paths, resembling a collection of directed star graphs [22]. As further detailed in the Supplementary Note 4, the normalized trace $\text{Tr}(W_{c})/N^{2}$ and the unnormalized $\text{Tr}(W_{c})$, both evaluated using the adjacency matrices $\tilde{A}$, in place of Jacobians, reinforce this observation by clustering gene regulatory networks closer to food webs and pathways, underscoring the role of perfect directedness in enhancing the passing property. We will explore these structural mechanisms further in the next section.
FIG. 4: Normalized trace of Gramian vs. the non-normality index. The normalized Henrici index $\hat{d_{F}}(A)$ of the Jacobian matrix $A$ is defined in Eq. (12). The data is displayed using the same color and marker conventions as those in Fig. 3. Pathway, Food web, Genetic, Connectome, and Power Grid networks are in black, red, yellow, purple, and green color, respectively.